# Modeling Binary Lenses and Sources with the BAGLE Python Package

### v1

@@ 7 Validation of Models
Lastly, a summary of binary modeling capabilities between different packages is provided in Table 3. pyLIMA is included in this summary. All four packages support static binary lenses and sources. In terms of orbital motion of binary lenses, all packages can simulate full Keplerian motion or circular “approximations” that are close to the actual solutions. BAGLE can model linear and accelerated approximations, which are computationally resourceful in microlensing events with very long orbital periods. MulensModel can also support simpler approximations with the inputs $\frac{ds}{dt}$ (rate of change of binary lens separation) and $\frac{d\alpha}{dt}$ (rate of change of the angle between the binary axis and the proper motion vector). In terms of the magnification algorithm used for binary lenses, BAGLE uses a polynomial root solver; VBMicrolensing provides multiple methods that include the polynomial root sovler, the Newton-Raphson method or contour integration; MulensModel provides an adaptive contouring method (M. Dominik, 2007), a hexadecapole/quadrupole approximation (uses Taylor expansion to evaluate point-source magnification at 9 and 13 points respectively) or the option to use VBMicrolensing as the backend; pyLIMA only uses VBMicrolensing as the backend. In terms of orbital motion for binary sources (often referred to as “xallarap”), all packages support circular orbital motion. MulensModel can handle elliptical orbital motion and BAGLE provides a linear and accelerated approximation too along with full elliptical orbital motion. It is important to re-iterate that pyLIMA and MulensModel implement a photometry-only fitting using their binary lens and source models. RTModel (developed to use VBMicrolensing for fitting events) can support joint photometric and astrometric fitting for all binary models except BSBL. BAGLE can support joint photometric and astrometric fitting of all binary models (i.e., PSPL, PSBL, BSPL, BSBL).
8 Results
@@
  • §8.2 - Centroid Shift Maps.
  • §8.3 - Dependency of mass ratio on caustic structures.
  • §8.4 - Dependency of orbital motion on fitting.
8.1 Results: Magnification Maps
Figure 11: Magnification map for a BSBL microlensing event with both the source and lens at $t_{p}$. The contour of the magnification map is indicative of the caustic. The color scale indicates ray density.
By solving the lens equation, the magnification of the source at any (projected) position relative to the position of the lens can be calculated. The magnification map is a visualization reflecting the calculated magnification at a given point. Magnification maps are generated using the inverse ray shooting method (D. P. Bennett, 2010). The inverse ray shooting method is a simple way to invert the lens equation by shooting rays backwards from the observer to the lens.
  • Create a sample grid of image positions in the lens frame of reference.
  • Calculate information about the lens, i.e., the lens positions at the time of closest approach (between the source and the lens system’s center of mass) and the lens mass ratio.
  • Use the lens mass ratio, lens positions, the image grid, and the source position to get the lensed source positions.
  • By mapping the image grid to the source plane and then binning the resulting source positions, the magnification map indicates how many rays fall into each pixel in the source plane grid. More rays imply a higher magnification. The magnification map we present in Figure  11 is purely statistical.
The magnification map for a BSBL event at $t_{p}$ is presented in Figure  11. The large boundaries of amplification in the magnification map are caustics. Caustics are characteristic features of the photometric curves of binary lenses. Since the sources have a finite size, their amplification does not become infinitely high. However, it still presents itself in the form of sharp peaks in our lightcurves.
A magnification map can be generated using a call to the get_magnification_map function in plot_models.py. The function performs the inverse-ray-shooting method and plots a magnification map at $t_{p}$ by default. An arbitrary time can be provided as an argument when calling the function.
8.2 Centroid Shift Maps
Figure 12: A color map for the centroid shift during a PSBL microlensing event with the lenses at $t_{p}$. The color scale indicates the absolute maximum centroid shift ($\boldsymbol{\delta_{c,obs}}$) in the image plane.
Like the magnification map, a forward ray shooting method can be used to shoot from the source plane to the image plane, creating a color map of the absolute maximum centroid shift. This forward ray shooting method works in the following way:
@@ 8.2 Centroid Shift Maps
A centroid shift color map for a PSBL event at $t_{p}$ is presented in Figure 12. A centroid shift map can be generated using a call to the get_centroid_shift_map function in plot_models.py. The function performs the forward-ray-shooting method and plots a magnification map at $t_{p}$. Like magnification maps, an arbitrary time can be provided as an argument when calling the function.
8.3 Results: Dependency of Mass Ratio and Separation
Figure 13: Caustics (pink) and critical curves (blue) for a PSBL model with the lenses at a separation of 5 mas. We present panels with $q$ ranging from 0.1 to 1.
Figure 14: Caustics (pink) and critical curves (blue) for a PSBL model at $t_{p}$ with a fixed $q=1$. The value of $a$ (provided as an input to the PSBL models) varies between 33 mas, 66 mas, 99 mas in the three panels.
In this section, we explore the dependency of caustics on mass ratios and separation. The separation is fixed at an arbitrarily chosen value of 3 mas and the mass ratio $q=\frac{m_{L,s}}{m_{L,p}}$ for a dark, non-planetary lens, ranges from 0.1 to 1. The caustics and critical curves are presented in Figure 13. We see that the caustic becomes more symmetric as $q\rightarrow 1$. The degree of asymmetry is directly dependent on $q$.
Similarly, by fixing the mass $q=\frac{m_{L,s}}{m_{L,p}}=1$, BAGLE can replicate a dark, non-planetary lens at varying angular separations between the primary and secondary lens. From the caustics and critical curves presented in Figure 14, we see that the caustic remains symmetric regardless of separation for a fixed $q=1$. At larger separations, the caustics also become increasingly elongated or stretched. Detailed descriptions of the geometry of caustics for the q = 1 case can be found in P. Schneider & A. Weiss (1986).
@@ 8.3 Results: Dependency of Mass Ratio and Separation
8.4 Results: Orbital Motion
@@
In our fitting process, BSPL models with linear, circular, and elliptical orbital motion are utilized to demonstrate how incorporating orbital motion enhances the quality of the fitting process. We present the photometric fitting results in Figure 15, the astrometric RA fitting results in Figure 16 and the astrometric Dec fitting results in Figure 17.
@@
Furthermore, the reduced chi-squared ($\bar{\chi}^{2}$) values are summarized in Table 5, and capture the quality of the fits. In our reduced chi-squared test, we calculated the degrees of freedom by subtracting the number of fitting parameters from the total number of astrometric and photometric data points. The best-fit model with linear orbital motion has a $\bar{\chi}^{2}=3.2$. The linear model overestimates the analytical uncertainties on the dataset. On the other hand, the best fits with circular and elliptical orbital motion have $\bar{\chi}^{2}=1.3$ and $\bar{\chi}^{2}=1.0$; these values indicate that the models with circular and elliptical orbital motion significantly improve our fitting results, and the residual difference between observed and fitted data is almost consistent with the error variance for the elliptical orbit model.
@@
From our reduced chi-squared test, we conclude that incorporating Keplerian orbital motion into BAGLE is necessary to create best-fit models with good fitting for complex lightcurves.
9 Conclusion
@@
From our fitting procedure using a mock dataset that replicates a binary-source, point-lens event, we conclude that the inclusion of orbital motion in binary microlensing events helps model complex photometric light curves. In these simulations, the accuracy of our binary fits based on $\bar{\chi^{2}}$ values improves with orbital motion.
@@
BAGLE’s capabilities for handling point-source, point-lens events are presented in J. Lu et al. (submitted), where BAGLE was compared with other microlensing packages like VBMicrolensing, pyLIMA, and MulensModel in detail. This paper includes a brief comparison between the different microlensing packages for point-source, binary-lens and binary-source, point-lens events. For point-source, binary-lens events, the residual difference in amplification between VBMicrolensing and BAGLE ranged around $10^{-2}$, as opposed to MulensModel. For binary-source, point-lens events, the residual difference in amplification ranged around $10^{-2}$ between all three packages.
@@
BAGLE’s model-fitting runtimes are comparable to VBMicrolensing and MulensModel for pre-instantiated binary-source, point-lens events. A large-scale simulation of multiple events is fastest with BAGLE for binary-source, point-lens events. BAGLE’s dependence on a polynomial root solver affects its runtime (full and pre-instantiated) for point-source, binary-lens calculations. Our future work involves exploring ways to improve the efficiency of the root solver through the use of JAX. We also aim to develop a similar lightcurve comparison for binary-source, binary-lens models, and compare events with orbital motion.
@@
In conclusion, the wide array of models and parameterizations available in BAGLE make it suitable for a joint photometric and astrometric fitting of binary events, including binary-source, binary-lens (BSBL) events. BAGLE’s new binary models will be used to work with data from the Vera C. Rubin Observatory, the Nancy Grace Roman Telescope, and other surveys. It will be used to better characterize measured microlensing signals of black hole astrometric candidates. These new models, which accurately capture the orbital dynamics of binary systems, will enhance our search for dark lenses, such as black holes, exoplanets, free-floating planets, and other intriguing candidates.
@@
We thank David Bennett, Valerio Bozza, Etienne Bachelet, and Radek Poleski for comments on package comparisons. We thank the Roman Galactic Exoplanet Project Infrastructure Team (RGES-PIT) working groups on microlens modeling and astrometry for useful discussions. The authors acknowledge support from the National Science Foundation under grant No. 2108185, the Heising-Simons Foundation under grant No. 2022-3542, and the Association of Universities for Research in Astronomy Space Telescope Science Institute under grant No. HST-GO-17081.004-A and HST-GO-16658.001-A. In Appendix A we describe the Thiele-Innes constants. In Appendices B and C, we present the coordinate transformation in binary microlensing to an arbitrary point along the binary. For example, transforming from a binary lens with respect to the primary to a binary lens with respect to the center of mass of the binary. This changes the measured closest approach distance and time which leads to a nontrivial transformation. These transformations can be used for both binary lens and binary source and can transform to any point along the binary axis. In Appendix  B, we transform $u_{0}$ and in Appendix  C, we transform $t_{0}$.
Appendix A Finding Thiele-Innes Constants
@@
The eccentric anomaly (E(t)) can be found using the mean anomaly and the eccentricity of the orbit as follows:
@@
Next, we find the Thiele-Innes Constants for the Keplerian orbits. These constants are solely used to transform the Keplerian orbital parameters into a partially linear basis, making it easier to find the binary system’s trajectory.
The Thiele-Innes Constants for the primary object are:
@@
$H_{pri}$ $=$ $\aleph_{pri}\left(\cos\omega_{pri}\sin\textit{i}\right)$ (A6)
In how we define our orbital parameterization, the only things we vary between the primary and secondary celestial objects are the length of the semi-major axis ($\aleph_{pri}$ and $\aleph_{sec}$) and the argument of periastron ($\omega_{sec}=\omega_{pri}+180\degree$). Therefore, the Thiele-Innes Constants for the secondary object are
@@
$H_{sec}$ $=$ $\aleph_{sec}\left(\cos\omega_{sec}\sin\textit{i}\right)$ (A7)
Appendix B $u_{0}$ transformation
We can think of going from one $u_{0}$ to another as a coordinate transformation from one point along the binary axis to another point (see Figure  18). We can transform from the geometric midpoint to the center of mass, to the primary, or to any other point. $L$ is the initial position on the binary axis and $L^{\prime}$ is the final position on the binary axis. $S$ is the closest the source gets to $L$ which occurs at time $t_{0}$ and $S^{\prime}$ is the closest the source gets to $L^{\prime}$ which occurs at time $t_{0}^{\prime}$. The source is moving with a velocity $\boldsymbol{\mu_{rel}}$ (in the frame of the lens). We define a coordinate system $R$ with $L$ at the center:
$L$ $=$ $[0,0]$ (B1)
$S$ $=$ $[u_{0,E},u_{0,N}]$ (B2)
$L^{\prime}$ $=$ $[d\hat{s}_{E},d\hat{s}_{N}]$ (B3)
$S^{\prime}$ $=$ $S+\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})=[u_{0,E}+\frac{\mu_{rel,E}}{\theta_{E}}(t_{0}^{\prime}-t_{0}),u_{0,N}+\frac{\mu_{rel,N}}{\theta_{E}}(t_{0}^{\prime}-t_{0})]$ (B4)
where $\boldsymbol{d}$ is the distance in units of $\theta_{E}$ along the binary axis that we transform. We normalize the by $\theta_{E}$ since $\boldsymbol{u_{0}}$ is in units of $\theta_{E}$.
We then define another coordinate system $R^{\prime}$ with $L^{\prime}$ at the center:
$L$ $=$ $[-d\hat{s}_{E},-d\hat{s}_{N}]$ (B5)
$S$ $=$ $S^{\prime}+\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}-t_{0}^{\prime})=[u_{0,E}+\frac{\mu_{rel,E}}{\theta_{E}}(t_{0}-t_{0}^{\prime}),u_{0,N}+\frac{\mu_{rel,N}}{\theta_{E}}(t_{0}-t_{0}^{\prime})]$ (B6)
$L^{\prime}$ $=$ $[0,0]$ (B7)
$S^{\prime}$ $=$ $[u_{0,E}^{\prime},u_{0,N}^{\prime}]$ (B8)
So when we transform from $R\rightarrow R^{\prime}$ we subtract $d\hat{s}$ since we shift the center from $[0,0]$ to $d\hat{s}_{E},d\hat{s}_{N}]$:
$L^{R^{\prime}}$ $=$ $L^{R}-d\hat{s}$ (B9)
$S^{R^{\prime}}$ $=$ $S^{R}-d\hat{s}$ (B10)
$L^{\prime R^{\prime}}$ $=$ $L^{\prime R}-d\hat{s}$ (B11)
$S^{\prime R^{\prime}}$ $=$ $S^{\prime R}-d\hat{s}$ (B12)

### v3

@@ VII Validation of Models
Lastly, a summary of binary modeling capabilities between different packages is provided in Table 5. pyLIMA is included in this summary. All four packages support static binary lenses and sources. In terms of orbital motion of binary lenses, all packages can simulate elliptical or circular Keplerian orbits that are close to the actual solutions. BAGLE can model linear and accelerated approximations, which are computationally resourceful in microlensing events with very long orbital periods. MulensModel also supports simpler approximations with the inputs $\frac{ds}{dt}$ (rate of change of binary lens separation) and $\frac{d\alpha}{dt}$ (rate of change of the angle between the binary axis and the proper motion vector). In terms of orbital motion for binary sources (often referred to as “xallarap”), all packages support circular orbital motion. MulensModel can handle elliptical orbital motion and BAGLE provides a linear and accelerated approximation too along with full elliptical orbital motion. It is important to re-iterate that pyLIMA and MulensModel implement a photometry-only fitting using their binary lens and source models. RTModel (developed to use VBMicrolensing for fitting events) can support joint photometric and astrometric fitting for all binary models (including finite sources) except binary-source, binary-lens models. BAGLE can support joint photometric and astrometric fitting of all binary models without finite sources. Note that BAGLE still provides support for finite source effects with point lenses.
VIII Results
@@
  • §VIII.2 - Centroid Shift Maps.
  • §VIII.3 - Dependency of Caustics on Mass Ratio and Separation.
  • §VIII.4 - Dependency of Fitting on Orbital Motion.
VIII.1 Results: Magnification Maps
Figure 11: Magnification map for a BSBL microlensing event with both the source and lens. Regions higher on the color scale indicate regions with higher magnification. The magnification map produced in BAGLE for a BSBL event only depends on the lens geometry. The secondary source in the map contributes through flux-weighting. Therefore, this magnification map would be the same for an equivalent PSBL event with the same lens geometry.
By solving the lens equation, the magnification of the source at any (projected) position relative to the position of the lens can be calculated. The magnification map is a visualization reflecting the calculated magnification at a given point. Magnification maps are generated using the inverse ray shooting method D. P. Bennett (2010), which inverts the lens equation by shooting rays backward from the observer to the lens plane.
To make a magnification map in BAGLE, these are the steps we follow:
  • Create a grid of image positions in the lens frame of reference.
  • By mapping the image grid to the source plane and then binning the resulting source positions, we produce a map of ray density, i.e., a map of the number of rays that land at each source position.
  • The source plane is pixelated, and the number of rays per pixel is proportional to the magnification of a source located at that pixel. The ray density map is normalized by the mean unlensed ray density to produce a magnification map.
The magnification map for a BSBL event is presented in Figure  11. The event has a lens mass ratio $q=1$ and a separation of $6$ mas between the lens at $t_{p}$. The Einstein radius is $4.9$ mas. At $t_{p}$, the sources have a separation of $13$ mas. The primary source has a magnitude of $16$, the secondary source has a magnitude of $20$ and $b_{sff}=1$. Note that the secondary source contributes through flux-weighting, which is why a singular caustic structure is visible on the map in Figure 11.
A magnification map can be generated using a call to the get_magnification_map function in plot_models.py. The function is called on a model instance, where the complete source and lens trajectory are specified. The get_magnification_map function can also be used on a PSBL model instance.
VIII.2 Centroid Shift Maps
Figure 12: A color map for the centroid shift during a PSBL microlensing event. The color scale indicates the absolute centroid shift ($\boldsymbol{\delta_{c,obs}}$) in the image plane.
Like the magnification map, a forward ray shooting method can be used to shoot from the source plane to the image plane, creating a color map of the absolute centroid shift. This forward ray shooting method works in the following way:
@@ VIII.2 Centroid Shift Maps
A centroid shift color map for a PSBL event is presented in Figure 12. The event has a lens mass ratio $q=1$ and a separation of $5$ mas between the lenses at $t_{p}$. The Einstein radius is $5.2$ mas. The source has a magnitude of 20 and $b_{sff}=1$.
A centroid shift map can be generated using a call to the get_centroid_shift_map function in plot_models.py. The function performs the forward-ray-shooting method and plots a centroid shift map. Like magnification maps, an arbitrary time can be provided as an argument when calling the function.
VIII.3 Results: Dependency of Caustics on Mass Ratio and Separation
Figure 13: Caustics (pink) and critical curves (blue) for a PSBL model with the lenses at a separation of 5 mas with $\theta_{E}=6$ mas. We present panels with $q$ ranging from 0.1 to 1.
Figure 14: Caustics (pink) and critical curves (blue) for a PSBL model at $t_{p}$ with a fixed $q=1$ with $\theta_{E}=6$ mas. The value of $a$ (provided as an input to the PSBL models) varies between 33 mas, 66 mas, 99 mas in the three panels.
In this section, we explore the dependency of caustics on mass ratios and separation. The separation is fixed at an arbitrarily chosen value of 5 mas and the mass ratio $q=\frac{m_{L,s}}{m_{L,p}}$ for a dark, non-planetary lens, ranges from 0.1 to 1. The Einstein radius is 6 mas. The caustics and critical curves are presented in Figure 13. We see that the caustic becomes more symmetric as $q\rightarrow 1$. The degree of asymmetry is directly dependent on $q$ (K. Griest & N. Safizadeh, 1998).
Similarly, by fixing the mass $q=\frac{m_{L,s}}{m_{L,p}}=1$, BAGLE can simulate a dark, non-planetary lens at varying angular separations between the primary and secondary lens. The Einstein radius is again 6 mas. From the caustics and critical curves presented in Figure 14, we see that the caustic remains symmetric regardless of separation for a fixed $q=1$. At larger separations, the caustics also become increasingly elongated or stretched. Detailed descriptions of the geometry of caustics for the q = 1 case can be found in P. Schneider & A. Weiss (1986).
@@ VIII.3 Results: Dependency of Caustics on Mass Ratio and Separation
VIII.4 Results: Dependency of Fitting on Orbital Motion
@@
In our fit, BSPL models with linear, circular, and elliptical orbital motion are utilized to demonstrate how incorporating orbital motion enhances the quality of the fit. We present the photometric fitting results in Figure 15, the astrometric RA fitting results in Figure 16 and the astrometric Dec fitting results in Figure 17.
@@
Furthermore, the reduced chi-squared ($\bar{\chi}^{2}$) values are summarized in Table 7, and capture the quality of the fits. In our reduced chi-squared test, we calculated the degrees of freedom by subtracting the number of fitting parameters from the total number of astrometric and photometric data points. The best-fit model with linear orbital motion has a $\bar{\chi}^{2}=3.2$. On the other hand, the best fits with circular and elliptical orbital motion have $\bar{\chi}^{2}=1.3$ and $\bar{\chi}^{2}=1.0$; these values indicate that the models with circular and elliptical orbital motion significantly improve our fitting results, and the residual difference between observed and fitted data is almost consistent with the error variance for the elliptical orbit model.
@@
From our reduced chi-squared test in Table 7, we conclude that incorporating Keplerian orbital motion into BAGLE is necessary to create best-fit models that fit well for complex lightcurves.
IX Conclusion
@@
From our fitting procedure using a mock dataset that replicates a binary-source, point-lens event, we conclude that the inclusion of orbital motion in binary microlensing events helps model complex photometric lightcurves. In these mock fitting, the accuracy of our binary fits based on $\bar{\chi^{2}}$ values improves with orbital motion.
@@
BAGLE’s capabilities for handling point-source, point-lens and finite-source, point-lens events are presented in J. Lu et al. (submitted), where BAGLE was compared with other microlensing packages like VBMicrolensing, pyLIMA, and MulensModel in detail. This paper includes a brief comparison between the different microlensing packages for point-source, binary-lens and binary-source, point-lens events. For point-source, binary-lens events and binary-source, point-lens events, the residual difference in magnification between VBMicrolensing and BAGLE ranged around $10^{-4}$. Between MulensModel and BAGLE, the difference ranged around $10^{-11}$ for point-source, binary-lens events and $10^{-15}$ for binary-source, point-lens events.
@@
BAGLE’s model-fitting runtimes are comparable to VBMicrolensing and MulensModel for pre-instantiated binary-source, point-lens events. A large-scale simulation of multiple events is fastest with BAGLE for binary-source, point-lens events. BAGLE’s dependence on a polynomial root solver affects its runtime (full and pre-instantiated) for point-source, binary-lens calculations. Our future work involves exploring ways to improve the efficiency of the root solver through the use of JAX. We also aim to develop a similar lightcurve comparison for binary-source, binary-lens models, and compare events with the inclusion of orbital motion.
@@
In conclusion, the wide array of models and parameterizations available in BAGLE make it suitable for a joint photometric and astrometric fitting of binary events, including binary-source, binary-lens events. BAGLE’s new binary models will be used to work with data from the Vera C. Rubin Observatory, the Nancy Grace Roman Telescope, and other surveys. These new models will be used to better characterize measured microlensing signals of black hole astrometric candidates. These new models, which accurately capture the orbital dynamics of binary systems, will enhance our search for dark lenses, such as black holes, exoplanets and other intriguing candidates.
@@
We thank David Bennett, Valerio Bozza, Etienne Bachelet, and Radek Poleski for comments on package comparisons. We thank the Roman Galactic Exoplanet Project Infrastructure Team (RGES-PIT) working groups on microlens modeling and astrometry for useful discussions. The authors acknowledge support from the National Science Foundation under grant No. 2108185, the Heising-Simons Foundation under grant No. 2022-3542, and the Association of Universities for Research in Astronomy Space Telescope Science Institute under grant No. HST-GO-17081.004-A and HST-GO-16658.001-A.
In Appendix A, we describe the method used to find the rectangular coordinates. In Appendix B we describe the Thiele-Innes constants. In Appendices C and D, we present the coordinate transformation in binary microlensing for $u_{0}$ and $t_{0}$. For example, BAGLE support binary lens parameterizations where the input parameters are in a reference frame with either the primary (or more massive) lens or the center of mass at origin. The different frames change the closest approach distance and time and the conversion between the two frames is non-trivial. These transformations can be used for both binary lens and binary source and can transform to any point along the binary axis. In Appendix  C, we transform $u_{0}$ and in Appendix  D, we transform $t_{0}$.
Appendix A Finding the rectangular coordinates
@@
The eccentric anomaly (E(t)) can be found using the mean anomaly and the eccentricity of the orbit by solving:
@@
Appendix B Finding Thiele-Innes Constants
We can find the Thiele-Innes Constants using Keplerian elements. The Thiele-Innes Constants for the primary object are:
@@ Appendix B Finding Thiele-Innes Constants
$H_{pri}$ $=$ $\aleph_{pri}\left(\cos\omega_{pri}\sin\textit{i}\right)$ (B1)
In how we define our orbital parameterization, the only parameters we vary between the primary and secondary celestial objects are the length of the semi-major axis ($\aleph_{pri}$ and $\aleph_{sec}$) and the argument of periastron ($\omega_{sec}=\omega_{pri}+180\degree$). Therefore, the Thiele-Innes Constants for the secondary object are
@@
$H_{sec}$ $=$ $\aleph_{sec}\left(\cos\omega_{sec}\sin\textit{i}\right)$ (B2)
Appendix C $u_{0}$ transformation
Transforming $u_{0}$ can be treated as a coordinate transformation of the origin from one point along the binary axis to another point (see Figure  18). We can transform the origin of the coordinate system from the geometric midpoint to the center of mass, to the primary, or to any other point. We preform the transformation in the heliocentric frame, so parallax does not need to be explicitly taken into account. Note that this coordinate transformation does not take orbital motion of the lens or source into account. It assumes that the angle of the binary axis and North and the direction of relative proper motion of the lens and source stay constant. $L$ is the initial position of the origin on the binary axis and $L^{\prime}$ is the final position of the origin on the binary axis. $S$ is the closest the source gets to $L$ which occurs at time $t_{0}$ and $S^{\prime}$ is the closest the source gets to $L^{\prime}$ which occurs at time $t_{0}^{\prime}$. The source is moving with a velocity $\boldsymbol{\mu_{rel}}$ (in the frame of the lens). We define a coordinate system $R$ with $L$ at the origin:
$L$ $=$ $[0,0]$ (C1)
$S$ $=$ $[u_{0,E},u_{0,N}]$ (C2)
$L^{\prime}$ $=$ $[d\hat{s}_{E},d\hat{s}_{N}]$ (C3)
$S^{\prime}$ $=$ $S+\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}^{\prime}-t_{0})=[u_{0,E}+\frac{\mu_{rel,E}}{\theta_{E}}(t_{0}^{\prime}-t_{0}),u_{0,N}+\frac{\mu_{rel,N}}{\theta_{E}}(t_{0}^{\prime}-t_{0})]$ (C4)
where $\boldsymbol{d}$ is the vector in units of $\theta_{E}$ along the binary axis that we transform. The magnitude of the vector is $d$ and the direction is $\hat{s}$ where $\hat{s}_{E}$ and $\hat{s}_{N}$ are the East and North components. We normalize the proper motions by $\theta_{E}$ since $\boldsymbol{u_{0}}$ is in units of $\theta_{E}$.
We then define another coordinate system $R^{\prime}$ with $L^{\prime}$ at the origin:
$L$ $=$ $[-d\hat{s}_{E},-d\hat{s}_{N}]$ (C5)
$S$ $=$ $S^{\prime}+\frac{\boldsymbol{\mu_{rel}}}{\theta_{E}}(t_{0}-t_{0}^{\prime})=[u_{0,E}+\frac{\mu_{rel,E}}{\theta_{E}}(t_{0}-t_{0}^{\prime}),u_{0,N}+\frac{\mu_{rel,N}}{\theta_{E}}(t_{0}-t_{0}^{\prime})]$ (C6)
$L^{\prime}$ $=$ $[0,0]$ (C7)
$S^{\prime}$ $=$ $[u_{0,E}^{\prime},u_{0,N}^{\prime}]$ (C8)
So when we transform from $R\rightarrow R^{\prime}$ we subtract $d\hat{s}$ since we shift the origin from $[0,0]$ to $d\hat{s}_{E},d\hat{s}_{N}]$:
$L^{R^{\prime}}$ $=$ $L^{R}-d\hat{s}$ (C9)
$S^{R^{\prime}}$ $=$ $S^{R}-d\hat{s}$ (C10)
$L^{\prime R^{\prime}}$ $=$ $L^{\prime R}-d\hat{s}$ (C11)
$S^{\prime R^{\prime}}$ $=$ $S^{\prime R}-d\hat{s}$ (C12)